Student population of a university has 30% asian, 40% american, 20% european and 10% african students. It is known that 40% of all asian students, 50% of all american students, 60% of all european students and 20% of all african students are girls. Find the probability that a girl chosen at random from the university is an asian?
1
Expert's answer
2020-06-21T16:49:28-0400

Let "A_1" denote asian student, "A_2" denote american student, "A_3" denote european student, "A_4" denote african student, and "G" denote that a tudent is a girl.

Given

"P(A_1)=0.3,"

"P(A_2)=0.4,"

"P(A_3)=0.2,"

"P(A_4)=0.1,"

"P(G|A_1)=0.4,"

"P(G|A_2)=0.5,"

"P(G|A_3)=0.6,"

"P(G|A_4)=0.2"


"P(G)=P(A_1)P(G|A_1)+P(A_2)P(G|A_2)+"

"+P(A_3)P(G|A_3)+P(A_4)P(G|A_4)="

"=0.3(0.4)+0.4(0.5)+0.2(0.6)+0.1(0.2)=0.46"

Use the Bayes' Theorem


"P(A_1|G)=\dfrac{P(A_1)P(G|A_1)}{P(G)}=""=\dfrac{0.3(0.4)}{0.46}={6\over 23}\approx0.260870"